# Elastohydrodynamics of 3D chemically active filaments

### v1

@@ 3.1.4 Diffusive-like motion: very high $E_{h}$
The pinwheeling trajectory, however, cannot become tighter indefinitely. Once $E_{h}\approx 34,000$, the pinwheeling regime gives way to ‘flap and tumble’ dynamics, as shown in Figure 6a. This regime is characterised by periods of coherent, periodic, antisymmetric flapping motion — where the filament makes reasonable forwards progress — punctuated by intervals of chaotic tumbling where the swimmer reorients. Viewed over longer timescales, this behaviour appears similar to that of diffusing particles.
For very high values of the elastohydrodynamic number, $E_{h}\gtrsim 50,000$, simulations show that, after some early-time transient planar motion, the chemically-active filament soon transitions to fully three-dimensional motion. The active forces are sufficiently strong compared to the elastic forces that the filament deforms significantly, resulting in sharp changes in direction and speed. This behaviour looks qualitatively like a random walk, and perhaps even chaotic. An example trajectory is shown in Figure 6b.
The progression in behaviour of the chemoelastic filaments can be clearly characterised by considering the midpoint displacement from its initial position over time, as shown in Figure 6c. As $E_{h}$ increases, the periodic, oscillatory motion at $E_{h}=30,000$ — corresponding to the stable periodic orbiting of Section 3.1.3 — is lost, with an initially more persistent motion at $E_{h}=40,000$ that stalls after some time, beyond which we observe displacement plots more characteristic of diffusion or random walks.
@@
Figure 6: Diffusion-like behaviour at high $E_{h}$. Snapshots of the trajectory of filament with (a) $E_{h}=37,000$ up to time $t=100$, and (b) $E_{h}=60,000$ up to a time $t=40$. After a short period of initial planar motion, the filament undergoes regular tumbling and reorientation in a chaotic, diffusion-like manner. (c) Filament displacement for various elastohydrodynamic numbers, with the legend denoting $E_{h}/1000$. As $E_{h}$ increases, we can see the progression from trapped periodic orbits, to more persistent directed motion, before reaching diffusive-like dynamics at very large $E_{h}$.
3.1.5 Persistent planar states
Looking at the early-time evolution from our planar S-shape initial condition, we see that near-planar trajectories that propel in a consistent direction with flapping back-and-forth can be observed prior to the above-mentioned later-time dynamics from around $E_{h}\approx 30,000$. For values beyond $E_{h}\approx 40,000$, these initial meandering trajectories persist for a noticeably longer time period. An example trajectory is shown in Figure 7a for $E_{h}=47,000$.
The progressive flapping motion comprises a travelling wave of curvature, that initiates in the middle of the filament, and alternates its direction of propagation towards either end of the filament, as seen in Figure 7b. As such, progress is a rocking back and forth motion along a straight trajectory.
Looking at the displacement over time for parameters $00040,000<E_{h}<50,$ 000, as shown in Figure 7c, we see that there is an intermediate band of elastohydrodynamic numbers in which this near-planar motion continues for long times. For values $00042,000\leq E_{h}\leq 47,$ 000, this motion persists for at least $t=100$. Additional simulations from a helical perturbation suggest that this near-planar dynamics can be reached from non-planar states, and so may be a stable attractor for some set of parameter values.
Outside of this band, the filament eventually deviates from these trajectories, showing that there are parameter values for which these dynamics are transient states, after which they follow the more diffusive-like motion observed for values of $E_{h}$ larger and smaller than this. For $E_{h}\gtrsim 50,000$, the filaments revert to diffusive behaviour and this persistent planar motion is lost.
@@ 3.1.5 Persistent planar states
Figure 7: Within a band of high values of $E_{h}$, a planar ‘flapping’ state exists. (a) Trajectory for $E_{h}=47,000$ for time $t=100$. (b) Curvature heatmap for the same simulation, showing waves of curvature travelling outwards from the centre of the filament. (c) The displacement over time for a range of values of $E_{h}$, showing persistent (planar) motion for intermediate $E_{h}$.
@@ 3.2.1 Planar translation and circling: low $E_{h}$
Figure 8: A stiff antisymmetrically-patterned filament, with $\mathcal{A}=\sin\pi s$, straightens and propels along its axis at low $E_{h}$. (a) The evolution of a filament for $E_{h}=4000$ up to time $t=10$. (b) The straight filament has slip flows directed towards the maximum in activity, and away from the minimum, resulting in the lower activity region towards the rear experiencing extension while the higher activity region at the front is under compression. If the active forcing is sufficiently high compared to the stiffness, then the region under compression may buckle.
When the filament is sufficiently stiff, elastic restoring forces dominate the dynamics, causing any shape perturbation to decay, and so the filament relaxes back to its straight configuration, propelling itself along its axis. We observe this behaviour for elastohydrodynamic numbers $E_{h}<5000$, with an example shown in Figure 8a.
The concentration gradients generated by the antisymmetric activity, $\mathcal{A}=\sin\pi s$, result in regions of both compression and extension in the filament, as shown in Figure 8b. The regions under compression are susceptible to buckling when the activity causes a strong enough flow relative to the material stiffness. Indeed, as $E_{h}$ increases further, beyond $E_{h}=5000$, the viscous forces acting on the self-propelled filament are sufficient to bend it. In particular, we observe deformation into a hook-shaped bend at the front end of the filament, with the corner of the hook centred close to the most active region (see Figure 9b).
Once in this configuration, there is a left-right asymmetry both in the distribution of the solute concentration, and hence slip velocity, due to the activity sink on the longer leg of the hook, as well as in the resulting drag on each of the legs due to their different lengths. The outcome is that the hook shape propels at an angle to the line that bisects the corner of the hook, resulting in an overall circling motion. An example of this is shown in Figure 9a for $E_{h}=6000$. As $E_{h}$ increases, the hook shape becomes more curved, and the trajectories follow a tighter circle.
@@
[…]
@@ 3.2.2 Transition to helical motion: moderate $E_{h}$
Figure 10: The filaments transition to three-dimensional motion via helical trajectories which become more pronounced with increasing $E_{h}$. (a) The helical pitch of trajectories as a function of $E_{h}$, with insets showing their helical motion. (b) Snapshots of a well-developed helical trajectory for $E_{h}=20,000$. These simulations were initiated with the filament in an initial helical shape.
@@
Similarly to the symmetric activity profiles, we observe a transition from planar motion to full three-dimensional dynamics at sufficiently high $E_{h}$. When this occurs, the planar circling behaviour appears to be metastable, as our initial planar perturbation settles into a fixed hook-shape circling behaviour on relatively short timescales, but a localised growth in torsion leads to out-of-plane motion. An example of the growth in torsion is shown in Figure 9c for $E_{h}=12,000$, demonstrating the exponential local growth of torsion. This localised torsion, typically observed between the midpoint of the filament and the most-curved region of the hook (positions highlighted in Figure 9b) for values of $E_{h}$ as low as 6000, destabilises the planar motion of the filament, twisting the filament out-of-plane into fully three-dimensional trajectories. However, for many of our simulations, we do not observe this transition because, although the torsion grows exponentially, it can be sufficiently slow that the torsion magnitude is not significant within the timeframes that we explore.
@@
To fully understand which values of the elastohydrodynamic number permit helical trajectories, we simulate the long-time dynamics of our chemoelastic filament when starting from an initial three-dimensional helical perturbation, and fit a helix to the resulting midpoint trajectory. In contrast to the sudden transition at a critical $E_{h}$ for the symmetrically-patterned filaments previously studied, this out-of-plane transition appears gradually as $E_{h}$ increases. The resulting pitch of this helix is plotted as a function of $E_{h}$ in Figure 10a, showing that, at lower $E_{h}$, there is a negligible pitch to the helix that may be difficult to observe except on extremely long timescales. As $E_{h}$ increases, the helix becomes more pronounced, with a larger pitch, allowing clear three-dimensional motion to be observed, with an example shown in Figure 10b.
Conversely, from the initial planar S-shape, the resulting helical trajectory only becomes apparent within a dimensionless time $t=100$ once the elastohydrodynamic number surpasses $E_{h}\approx 20,000$, although, as can be seen from Figure 10a, low-pitch helical paths exist for $E_{h}$ much lower than this. A typical example of this helical motion from the planar initial configuration is shown in Figure 11. The motion initially remains planar for some time before transitioning to a helical path. At larger $E_{h}$, the out-of-plane motion occurs sooner and the resulting helix pitch increases. During the helical motion, the base state of the filament is a hook shape with a localised twist out-of-plane. Around this, there is evidence of a travelling wave from the centre towards the tip from the heatmaps of curvature and torsion (Figure 11).
@@
Figure 11: The transition from planar to three-dimensional motion. Here, for $E_{h}=28,000$ up to time $t=30$. (a) The trajectory showing an initial planar circling, before a sharp transition towards a helical spiralling motion out-of-plane. Heatmaps of (b) curvature and (c) torsion, showing a clear transition from planar to non-planar motion. Note the rapid increase in torsion prior to out-of-plane dynamics.
3.2.3 Run-and-tumble and re-straightening: high $E_{h}$
When $E_{h}$ moves beyond approximately $40,000$, the helical dynamics are lost; the filament becomes more flexible leading to more chaotic-looking paths. Examples of these trajectories are given in Figure 12a,b. The resulting trajectories are reminiscent of a persistent random walk, or perhaps the run-and-tumble dynamics of bacteria, with intervals of relatively persistent trajectories with some lateral oscillations due to filament flapping, connected by sharp turns.
(a)(b)(c)
Figure 12: Beyond helical motion. Example trajectories up to time $t=50$, from left to right: $E_{h}=60,000$, $80,000$ and $90,000$. The dynamics deviate from helical trajectories at large $E_{h}$, showing run-and-tumble motion at high $E_{h}$, and eventual re-straightening of paths.
At very high $E_{h}>80,000$, the dynamic behaviour, perhaps surprisingly, becomes less chaotic, as shown in Figure 12c. Similar to the persistent meandering observed for the symmetric activity, we see flapping back and forth around a consistent trajectory, but here with some rotation about the direction of motion.
Both the run-and-tumble and straightening dynamical behaviour may be explained by looking more closely at the filament shape during the dynamics. As the filament propels forward, the front end rapidly buckles from side-to-side in a chaotic manner, while the rear portion of the filament remains relatively straight. This is due to the slip flow in the front half being locally compressional, leading to buckling, while the rear half experiences a locally extensional flow, straightening the filament out. Whereas at lower $E_{h}$, the filament was both being pulled forward by the front half and pushed by the rear, here the filament may be thought of as mostly being pushed from the back half, while the rapid oscillations of the front half cause small side-to-side deviations. In the case of run-and-tumble dynamics, these deviations may cause small random perturbations in the path direction, making the trajectory more like a persistent random walk; for the higher values of $E_{h}$, the front-end deformation may be so rapid and severe that their effect averages out over time.
@@ 3.3.1 A near constant, fore-aft symmetric activity profile
However, the values for $E_{h}$ where the behaviour transitions are noticeably higher. For instance, we first observe the stable periodic orbiting behaviours from $E_{h}\approx 20,000$, in contrast to the $17,000$ for $\mathcal{A}=\sqrt{1-s^{2}}$. This is perhaps because, for this near-constant activity, the slip flows are weaker over the majority of the filament since concentration gradients are reduced, and so either a higher propulsive force or a weaker bending rigidity is required to cause the appropriate buckling transitions.
@@ 3.3.2 Reversing the activity and mobility
In the examples considered so far, we have focused on chemical patterning with positive activity and mobility, i.e. the solute is generated at the filament surface and slip flows are directed up the surface concentration gradients. This choice resulted in straining flows that cause buckling in the symmetric filaments. A natural question to ask is: what happens if the activity and/or mobility change sign? Changing the sign of the activity causes the chemical patterning to deplete the solute, instead of generating it, and changing the sign of mobility causes slip flows to be directed down surface concentration gradients, rather than up them.
Due to the linearity of our system of equations, reversing the sign of both the activity and mobility results in the same slip flows and hence identical dynamics. This means that chemical patterning which both depletes the solute, $\mathcal{A}<0$, and causes slip flows down the chemical gradients, $\mathcal{M}<0$, have the same dynamics as those considered above. It is therefore only necessary to consider what happens when $\mathcal{A}<0$ and $\mathcal{M}>0$, which we consider for the chemical patterns from Sections 3.1 and 3.2.
@@
Meanwhile for any antisymmetric chemical patterning, there is no observed difference in dynamics; reversing the sign of the activity or mobility is equivalent to swapping the front and back of the filament, and so must behave in the same manner as previously discussed.
@@ 3.3.3 Turning off the azimuthal component
A key aspect of the theory of chemically-propelled filaments that is uncommon in other physical systems is that the azimuthal slip flow — acting around the slender cross-sections, perpendicular to the local tangent — is as important to the leading-order behaviour as the flow along the length of the filament, as can be seen in e.g. eqs. 5 and 7. It is not clear from our previous simulations what effect this component of the slip flow has on the observed dynamics.
To understand the importance of the azimuthal component of the flows, caused by the contribution of the ${c}^{(1)}$ component of the slip flow in Slender Phoretic Theory, as given in Equation 5, we ran simulations with this component set to zero. Qualitatively, we observe the same transitions in the dynamic behaviour for $\mathcal{A}=\sqrt{1-s^{2}}$. As $E_{h}$ increases, the dynamics progress from straightening to planar U-shape translation, followed by an out-of-plane transition, then pinwheeling orbits, and eventually paths that are more random and diffusive. However, these transitions occurred at significantly lower $E_{h}$ than the full simulations including the azimuthal component. For example, the initial buckling occurs below $E_{h}=1000$, with out-of-plane trajectories by $E_{h}=5000$ and pinwheeling at $E_{h}=6000$.
This perhaps suggests that the azimuthal component is stabilising the flow around the filament and resisting deformation to the next dynamic mode in the full-physics simulations. A possible explanation for this is that, when the filament is in a near-planar configuration, azimuthal components driven by curvature also lie close to the plane of the filament, and fluid travels equally around either side of the thin body, suppressing any out-of-plane symmetry-breaking instabilities unless the flow is much more dominant compared to the filament elasticity. When these components are removed, there are only tangential forces promoting buckling instabilities, meaning that this three-dimensional transition occurs for much lower values of $E_{h}$.

### v2

@@ 3.1.4 Diffusive-like motion: very high $E_{h}$
The pinwheeling trajectory, however, cannot become tighter indefinitely. Once $E_{h}\approx 49,000$, early-time transient dynamics gives way to complex, three-dimensional motion. The active forces are sufficiently strong compared to the elastic forces that the filament deforms significantly, resulting in sharp changes in direction and speed. This behaviour looks qualitatively like a random walk, and perhaps even chaotic. Example trajectories are shown in Figure 6a,b.
The progression in behaviour of the chemoelastic filaments can be clearly characterised by considering the midpoint displacement from its initial position over time, as shown in Figure 6c. The presence of significant ballistic-like motility can be seen clearly for $E_{h}=60,000$, with dynamics that are more similar to diffusion or random walks occurring for the other values shown.
@@
Figure 6: Diffusion-like behaviour at high $E_{h}$. Snapshots of the trajectory of filament with (a) $E_{h}=50,000$ up to time $t=40$, and (b) $E_{h}=70,000$ up to a time $t=40$. In each, time progresses from dark to light. After a short period of initial planar motion, the filament undergoes regular tumbling and reorientation in a chaotic, diffusion-like manner. (c) Filament displacement for various elastohydrodynamic numbers, with the legend denoting $E_{h}/1000$. As $E_{h}$ increases, we can see the progression from trapped periodic orbits, to more persistent directed motion, before reaching diffusive-like dynamics at very large $E_{h}$.
3.1.5 Persistent ballistic motion
Prompted by the large displacement seem in Figure 6c, we focus on the dynamics for values near $E_{h}=60,000$. as shown in Figure 7c. This reveals more examples of apparently meta-stable planar dynamics, characterised by sustained ‘flapping’ motion. An example trajectory is shown in Figure 7a for $E_{h}=60,000$.Additional simulations from a helical perturbation suggest that these dynamics can be reached from non-planar states, and so may be a stable (or metastable) attractor for some set of parameter values.
The progressive flapping motion comprises a travelling wave of curvature, that initiates in the middle of the filament, and alternates its direction of propagation towards either end of the filament, as seen in Figure 7b. As such, progress is a rocking back and forth motion along a nearly straight trajectory.
@@ 3.1.5 Persistent ballistic motion
Figure 7: Within a band of high values of $E_{h}$, a ballistic-like motion appears to persist. (a) Trajectory for $E_{h}=60,000$ for time $t=100$, with time progressing from dark to light. (b) Curvature heatmap for the same simulation, showing waves of curvature travelling outwards from the centre of the filament. (c) The displacement over time for a range of values of $E_{h}$, showing (planar) motion that persists for intermediate $E_{h}$.
@@ 3.2.1 Planar translation and circling: low $E_{h}$
Figure 8: A stiff, antisymmetrically-patterned filament, with $\mathcal{A}=\sin\pi s$, straightens and propels along its axis at low $E_{h}$. (a) The evolution of a filament for $E_{h}=2000$ up to time $t=2$, with time progressing from dark to light. (b) The straight filament has slip flows (denoted by arrows) directed towards the minimum in activity, and away from the maximum. The filament experiences viscous forces opposing the slip flow, resulting in the lower activity region towards the rear experiencing extension, while the higher activity region at the front is under compression. If the active forcing is sufficiently high compared to the stiffness, then the region under compression may buckle.
When the filament is sufficiently stiff, elastic restoring forces dominate the dynamics, causing any shape perturbation to decay. Hence, the filament relaxes back to its straight configuration, propelling itself along its axis. We observe this behaviour for elastohydrodynamic numbers approximately lower than $5000$, with an example shown in Figure 8a.
The concentration gradients generated by the antisymmetric activity, $\mathcal{A}=\sin\pi s$, result in regions of both compression and extension in the filament, due to the slip flows illustrated in Figure 8b. The regions under compression are susceptible to buckling when the activity causes a strong enough flow relative to the material stiffness. Indeed, as $E_{h}$ increases further, beyond $E_{h}=5000$, the viscous forces acting on the self-propelled filament are sufficient to bend it. In particular, we observe deformation into a hook-shaped bend at the front end of the filament, with the corner of the hook centred close to the most active region (see Figure 9b).
Once in this configuration, there is a left-right asymmetry both in the distribution of the solute concentration (and the associated slip velocity) due to the activity sink on the longer leg of the hook, as well as in the resulting drag on each of the legs due to their different lengths. The outcome is that the hook shape propels at an angle to the line that bisects the corner of the hook, resulting in an overall circling motion. An example of this type of motion is shown in Figure 9a for $E_{h}=28,000$. As $E_{h}$ increases, the hook shape becomes more curved, and the trajectories follow a tighter circle.
@@
Figure 9: Buckling of the filament results in a planar circling motion. (a) Example circling trajectory for $E_{h}=28,000$ for a time $t=30$ with time progressing from dark to light. (b) Steady state hook shapes for $E_{h}=6,000$$28,000$. Squares show the location of maximum torsion. (c) Torsion heatmap for $E_{h}=28,000$ showing a localised increase in torsion towards the filament centre. Inset: at a fixed point along the filament (given by the white line), the torsion increases exponentially over time.
@@ 3.2.2 Transition to helical motion: moderate $E_{h}$
Figure 10: Filaments transition to three-dimensional motion via helical trajectories that, up to a point, become more pronounced with increasing $E_{h}$. (a) The helical pitch of trajectories as a function of $E_{h}$, with insets showing their motion. Pitch is highly non-monotonic: a sharp maximum is visible near $E_{h}\approx 8,000$, with trajectories reducing their pitch and radius (not shown) after the peak. For the largest values of $E_{h}$ shown on this plot, the motion increasingly resembles quasi-planar dynamics, with slow out-of-plane drift. (b) Snapshots of a well-developed helical trajectory for $E_{h}=18,000$, with time progressing from dark to light. These simulations were initiated with the filament in a helical shape.
@@
Similarly to the symmetric activity profiles, we observe a transition from planar motion to full three-dimensional dynamics at sufficiently high $E_{h}$. Here, this transition occurs around $E_{h}=5500$, so that planar circling appears to only be stable for a small range of values. Commensurate with this, in this regime the planar circling behaviour appears to be metastable, as our initial planar perturbation settles into a fixed hook-shape circling behaviour on relatively short timescales, but a localised growth in torsion leads to out-of-plane motion on longer timescales. An example of the growth in torsion is shown in Figure 9c for $E_{h}=28,000$, demonstrating the exponential local growth of torsion. This localised torsion, typically observed between the midpoint of the filament and the most-curved region of the hook (positions highlighted in Figure 9b), destabilises the planar motion of the filament, twisting the filament out-of-plane into fully three-dimensional trajectories. Simulating the same motion for longer times reveals a seemingly robust helical trajectory.
@@
To fully understand which values of the elastohydrodynamic number permit helical trajectories, we simulate the long-time dynamics of our chemoelastic filament when starting from an initially three-dimensional helical perturbation, and fit a helix to the resulting midpoint trajectory. In contrast to the sudden transition at a critical $E_{h}$ for the symmetrically-patterned filaments shown in Figure 3, this out-of-plane transition appears gradually as $E_{h}$ increases. The resulting pitch of this helix is plotted as a function of $E_{h}$ in Figure 10a, showing that pitch initially grows with $E_{h}$. As $E_{h}$ increases, the helix becomes significantly more pronounced and emerges over lengthscales many times that of the filament. Curiously, after a peak around $E_{h}\approx 8,000$, the pitch decreases, leading to large, approximately circular trajectories that slowly move in three dimensions and trace out very large helices. Notably, these dynamics might easily be mistaken for planar motion on shorter timescales. An example of these dynamics, associated with significantly smaller scale helices than those before the peak in pitch, is shown in Figure 10b. We remark that the emergent radii of the helical trajectories follow the same trend as the measured pitches, such that the pitch serves as an approximate measure of the lengthscale of the helices.
These observations broadly align with simulations from an initially planar state, with helical motion of various forms present from around $E_{h}\approx 5,500$. This suggests that this helical behaviour is a stable attractor of the dynamics, though a full and proper treatment of the solutions to the steady state dynamics is warranted to formally address the notion of stability. A typical example of this form of helical motion from the planar initial configuration is shown in Figure 11.The large value of $E_{h}$ chosen here yields tight circles and a corresponding tight helix. The motion initially remains planar for some time before transitioning to a helical path. At larger $E_{h}$, the out-of-plane motion appears to occur sooner and the resulting helix pitch increases. During the helical motion, the base state of the filament is a hook shape with a localised twist out-of-plane. As visible in Figure 11b,c, the curvature of the filament remains approximately unchanged throughout the motion (after initial transients), whilst the torsion undergoes a significant change that corresponds to the filament leaving the plane.
@@
Figure 11: The transition from planar to three-dimensional motion. Here, we see helical motion for $E_{h}=50,000$ up to time $t=50$. (a) The trajectory showing an initial planar circling, before a sharp transition towards a helical spiralling motion out-of-plane, with time progressing from dark to light. Heatmaps of (b) curvature and (c) torsion, with only the latter showing a clear transition from planar to non-planar motion.
3.2.3 Run-and-tumble and beyond: high $E_{h}$
When $E_{h}$ moves beyond approximately $80,000$, the helical dynamics are lost; the filament becomes more flexible, leading to more chaotic-looking paths, beginning with irregular, tightly wound helices. Examples of these trajectories are given in Figure 12.The resulting trajectories are reminiscent of a persistent random walk, or perhaps the run-and-tumble dynamics of bacteria, with intervals of relatively persistent trajectories connected by sharp turns.
This run-and-tumble-like behaviour may be explained by looking more closely at the filament shape during the dynamics. As the filament propels forward, the front end rapidly buckles from side-to-side in a chaotic manner, while the rear portion of the filament remains relatively straight. This is due to the slip flow in the front half applying locally compressional forces, leading to buckling, while the rear half experiences a locally extensional forcing, straightening the filament out. Whereas at lower $E_{h}$ the filament was both being pulled forward by the front half and pushed by the rear, here the filament may be thought of as mostly being pushed from the back half, while the rapid oscillations of the front half cause small side-to-side deviations. These deviations may cause small, seemingly random perturbations in the path direction, making the trajectory more like a persistent random walk.
(a)(b)
Figure 12: Beyond helical motion. Example trajectories for (a) $E_{h}=80,000$ and (b) $90,000$, each up to time $t=50$, with time progressing from dark to light. The dynamics deviate from helical trajectories at large $E_{h}$, showing something akin to run-and-tumble motion at high $E_{h}$.
@@ 3.3.1 A near constant, fore-aft symmetric activity profile
However, the values for $E_{h}$ where the behaviour transitions are noticeably higher. For instance, we first observe the stable periodic orbiting behaviours from $E_{h}\approx 30,000$, in contrast to the $24,000$ for $\mathcal{A}=\sqrt{1-s^{2}}$. This is perhaps because, for this near-constant activity, the slip flows are weaker over the majority of the filament since concentration gradients are reduced, and so either a higher propulsive force or a weaker bending rigidity is required to cause the appropriate buckling transitions.
@@ 3.3.2 Reversing the activity and mobility
In the examples considered so far, we have focused on chemical patterning with positive activity and negative mobility, i.e. the solute is generated at the filament surface and slip flows are directed down the surface concentration gradients. This choice resulted in straining flows that cause buckling in the symmetric filaments. A natural question to ask is: what happens if the activity and/or mobility change sign? Changing the sign of the activity causes the chemical patterning to deplete the solute, instead of generating it, and changing the sign of mobility causes slip flows to be directed up surface concentration gradients, rather than up them.
Due to the linearity of our system of equations, reversing the sign of both the activity and mobility results in the same slip flows and, hence, identical dynamics. In other words, the sign of $\mathcal{A}\mathcal{M}$ drives the dynamics. This means that chemical patterning which both depletes the solute, $\mathcal{A}<0$, and causes slip flows up the chemical gradients, $\mathcal{M}>0$, have the same dynamics as those considered above. It is therefore only necessary to additionally consider what happens when $\mathcal{A}>0$ and $\mathcal{M}>0$, which we consider for the chemical patterns from Sections 3.1 and 3.2.
@@
Meanwhile, for antisymmetric chemical patternings, there is no observed difference in dynamics when reversing the sign of $\mathcal{A}\mathcal{M}$. Indeed, reversing the sign of only one of the activity or mobility is equivalent to swapping the front and back of the filament, and so yields unchanged dynamics.
@@ 3.3.3 Turning off the azimuthal component
A key aspect of the theory of chemically propelled filaments that is uncommon in other physical systems is that the azimuthal slip flow (acting around the slender cross-sections, perpendicular to the local tangent) is as important to the leading order behaviour as the flow along the length of the filament, as can be seen in e.g. eqs. 5 and 7. It is not clear from our previous simulations what effect this component of the slip flow has on the observed dynamics.
To understand the importance of the azimuthal component of the flows, caused by the contribution of the ${c}^{(1)}$ component of the slip flow in slender phoretic theory, as given in eq. 5, we ran simulations with this component artificially set to zero. Qualitatively, we observe the same transitions in the dynamic behaviour for $\mathcal{A}=\sqrt{1-s^{2}}$. As $E_{h}$ increases, the dynamics progress from straightening to planar U-shape translation, followed by an out-of-plane transition, then pinwheeling orbits, and eventually paths that appear more random and diffusive. However, these transitions occurred at significantly lower $E_{h}$ than the full simulations including the azimuthal component. For example, the initial buckling occurs below $E_{h}=1000$, with pinwheeling by $E_{h}=7000$.
This perhaps suggests that the azimuthal component is stabilising the flow around the filament and resisting deformation to the next dynamic mode in the full-physics simulations. A possible explanation for this is that the slip flow suppresses buckling. If the filament begins to curve, then the positive activity generates relatively higher solute concentration on the inside of the curve than outside; for the negative mobility considered, the slip flows are then directed from the inside of the bend to the outside, propelling the bend in the opposite direction, back towards a straight configuration (Makanga et al., 2025). When these azimuthal components are removed, there are only tangential forces promoting buckling instabilities, meaning that this three-dimensional transition occurs for much lower values of $E_{h}$.
We note that, in these simulations, the hydrodynamics are forced by effectively applying a distribution of tangential slip flows (generated by the chemical activity) that are independent of time and fixed in the body frame. Since we are using a linear local slender body theory for Stokes flow — resistive force theory — identical dynamics may be obtained by instead applying a distribution of tangential forces along the centreline, since tangential flow and force are linearly related. However, more generally, we do not expect this equivalence to hold for more accurate models, for example using non-local theories for chemical activity or flow.
Figure 13: Categorisation of the observed dynamic behaviours of the symmetric and antisymmetric chemical activities. In each case, as $E_{h}$ increases, we observe a planar buckling transition, followed by an out-of-plane transition. Beyond this we see an array of dynamics, including closed orbits, helical trajectories, near-planar states and chaotic diffusion-like motion.